# Is my Data in your AI Model? Membership Inference Test with Application to Face Images

### v2

@@ Abstract
I Introduction
In June 2023, the European Parliament adopted its position with respect to Artificial Intelligence (AI) [1]. Concretely, they have imposed some obligations to all AI companies to guarantee the protection of the citizen rights. For example, this new regulation imposes the registration of AI/ML models in a European Union (EU) database and gives national authorities of EU countries the power to request access to both the trained AI models and the associated procedural details conducted in the development of those models. This new regulation is a game changer, imposing more transparency in the development of AI technologies in Europe. As a result, novel auditing tools based on the access to trained AI models must be developed to monitor AI technologies and their correct deployment in our society.
@@ I Introduction
This objective is related to the field of Membership Inference Attacks (MIAs), which refers to privacy attacks that target AI models trained on sensitive data [14, 15]. MIAs seeks to infer membership information about the training dataset used by exploiting the behavior and responses of the model. Also, MIAs leverage the potential memorization of the training data used by the AI models, especially when the models overfit [16, 17] or when the training data contain unique or distinctive patterns [14]. However, the performance of MIAs has been questioned in the last years due to the inherent complexity of the task [18, 19].
@@
While MIAs are traditionally considered as attacks to quantify the potential privacy leakage of AI models, in the present article we introduce a novel perspective in line with the new AI regulation [1]. Concretely, we consider it as an auditing tool for detecting the potential use of unauthorized data for training AI models, a concept we term as Membership Inference Test (MINT). The aim of MIAs and MINT is similar: to determine whether a given data sample was used during the training of a target AI model. However, although MIAs and our proposed MINT share the same objective, they diverge in their environmental conditions, leading to different methodological approaches. In MIAs, the model is trained assuming the role of an attacker who does not have access to the original model, but to a replica as similar as possible to it (shadow model) [14]. In MINT, we assume the role of an auditor who has direct access to the original model or partial information of it (e.g., intermediate activations). This is aligned with the new AI regulation [1], which imposes the registration of the trained AI models in an EU database. As a result, this new paradigm moves from the attacker to the auditor point of view, allowing the development of MINT models under different environmental conditions.
@@
In the present study we concentrate on the challenging task of Face Recognition, due to the sensitive nature of face biometrics data [20, 21, 22] according to the recent General Data Protection Regulation (GDPR) promoted by EU [23]. Although MIAs and MINT differ due to their distinct environmental conditions, they also share significant similarities, so methods in one domain can inspire advancements in the other as well. To the best of our knowledge, there is no literature on the application of MIAs to Face Recognition. This gap in the state of the art is particularly crucial as MIAs frequently deal with classification tasks, using the network’s output vector to determine membership in the training data, like a form of ‘probability’ of belonging to each class [19, 24, 25, 26, 16]. However, due to the characteristics of Face Recognition systems, focused on recognition rather than classification, such vector usually does not exist. This multiclass probability vector is traditionally replaced by an embedding feature vector representing individual characteristics in a learned space. Consequently, the methods and results diverge even more from what we typically see in state-of-the-art MIAs studies [19].
@@
  • We introduce MINT as a novel perspective to detect if given data was used during the training process of AI models from an AI auditor standpoint.
@@
The article is structured as follows. Sec. II provides a revision of the state of the art in the scope of the article. Sec. III describes the main concepts of the proposed MINT, and the specific details for the task considered, Face Recognition. The databases, Face Recognition systems, and experimental protocol are described in Sec. IV. Results are discussed in Sec. V, whereas conclusions and future work are finally drawn in Sec. VII.
II Related Works
First, it is important to remark that, to the best of our knowledge, the specific problem addressed in the present article has not been directly tackled in the current literature. Nevertheless, there are studies and research directions that touch upon similar topics or are grounded in analogous principles, which we will briefly discuss in this section. Our study closely relates to three fundamental research directions: “Membership Inference Attacks”, “Reconstructing Training Data from Neural Networks”, and “Data Tracing”. In the following, we provide concise explanations of these three primary research areas, with special emphasis on MIAs due to its significant similarity to our own research.
II-A Membership Inference Attacks
The capability to detect the data used for training AI models presents a security concern that can potentially unveil sensitive [2] or private information [27], encompassing areas such as shopping preferences, health records, and pictures, among many others. TThis capability has been explored in the literature mainly under the name of Membership Inference Attacks (MIAs). These attacks aim to extract such sensitive information through an adversarial approach, without access to the model or its training data, as providers are hesitant to disclose such data. The pioneering work by Shokri et al. [14] laid the groundwork for this line of research. In their approach, they used “shadow models” that emulate the functionality of the original model, despite having no direct access to it. To construct these so-called “shadow models”, it is essential to have knowledge about the architecture of the original model they are based on, along with a subset of samples and statistics derived from the original dataset used for training. As a result, they crafted “shadow training sets” mirroring the original dataset, subsequently training shadow models to mimic the original model. This approach offers complete control over the training and non-training data of these models. A binary classifier was trained to distinguish between the training and non-training data used in the shadow models. Images were passed through the network, and output embeddings from all shadow models were used to train a simple binary classifier. Better results were observed with an increasing number of shadow models.
@@ II-A Membership Inference Attacks
More recently, the authors of [18] delved into the challenges of MIAs and the inherent complexity of the task. They concluded that the actual performance in the task is notably worse than the results published in the state of the art. This is largely due to the insufficient evaluation methodology. Furthermore, they explored previously untapped White-box information involving gradients, although they were not able to significantly improve the Black-box framework performance.
Finally, it is noteworthy to mention a series of studies that may look similar [32, 33, 34]. These studies aim to detect users who have been utilized in training a Face Recognition system through distances between users in the embedding space [32, 34] or via guided captioning [33]. While these studies are intriguing, they diverge significantly in terms of objectives and methodologies from both our proposed method and the broader MIAs and MINT research lines.
II-B Reconstructing Training Data from AI Models
@@ II-C Data Tracing
III Membership Inference Test (MINT): Proposed Approach
MINT is a Membership Inference Test for a neural network’s training data that aims to comply with current legislation [1], hence, assuming a certain level of collaboration from the model developer. The ultimate goal is to develop techniques for all data types (text, images, audio, etc.), taking into account different amounts of information provided by the model developer about the model training and operation. This spans from scenarios where no data information is provided (unsupervised) or very little (semi-supervised), up to when sufficient information about the training data is granted (supervised). Simultaneously, it considers whether you have access to the entire model (white-box) or merely to the model’s output (black-box). In the present study, we analyze a broad range of scenarios, varying the quantity of data available for training the MINT model and contemplating both white-box and black-box settings.
Detecting the data used to train a model is a complex task in some sense opposite to the primary goal of common machine learning (ML) problems: generalize well during inference to samples unseen during learning, i.e., behaving similarly for both training and non-training samples. Consequently, in standard ML we expect samples of the same class to be proximate in the model representation space, regardless of their presence in the training set (Fig. 2). MIAs and MINT on the other hand, seek to distinguish samples (used or not for training the Audited Model $M$) that the model $M$ intends to represent as close as possible. Hence, much of the existing MIA literature revolves around overfitted networks that do not generalize well [14, 47, 48, 19, 49, 50].
In MINT, we assume the role of model auditors with access to the original model, or at least to partial information of it. This is a consistent scenario either due to the legal considerations outlined in the present study (i.e., granting national authorities the possibility to request the model [1]), or as a result of developer transparency initiatives aimed at end-users. This is an important distinction from related works around MIAs, as here we do not need to train “shadow models” to simulate the original model’s behavior. The “shadow models”, while a useful approximation, operate in a space different from the original (Fig. 2 left), potentially imposing limitations on the achievable performance [18]. Instead, with access to the original model, we can directly apply our proposed approach, which may yield different detection results and methods for similar tasks.
@@ III Membership Inference Test (MINT): Proposed Approach
III-A Problem Statement and Terminology
Let us consider a Training Dataset ($\mathcal{D}$), an External Dataset ($\mathcal{E}$) and a collection of samples $d$ $(d\in\mathcal{D}\cup\mathcal{E})$. We assume a learned model ($M$) that is trained for a specific task (e.g., text generation, Face Recognition, etc.) using the dataset $\mathcal{D}$. For any input data record ($d$), the model ($M$) generates an outcome ($y$) based on $d$ and a set of parameters (w) learned during the training process, i.e., $y=M(d|\textbf{w})$.
@@ III-A Problem Statement and Terminology
  • External Data $\mathcal{E}$: any data out of the collection ($\mathcal{D}$).
@@
  • Auxiliary Auditable Data $AAD=N(d|\textbf{w}^{\prime})$: intermediate outcomes of $M$ that result from processing an input data $d$ using a subset $\textbf{w}^{\prime}$ of the parameters w. The model outcome can be seen as the case where $^{\prime}=\textbf{w}$ w in which case $N(d|\textbf{w}^{\prime})=M(d|\textbf{w})$.
@@
III-B MINT: Application to Face Recognition Models
In the previous Sec. III-A, we have introduced the general concepts of our MINT proposal, which can be applied to any type of Audited Model $M$. Next, we detail the various components of the proposal for the application of MINT to the challenging task of Face Recognition.
@@ III-B MINT: Application to Face Recognition Models
III-C Proposed MINT Models
@@
  • Vanilla MINT Model: We propose a MLP network to capture the training patterns within the vectorized information of the audited model ($M$). When it comes to the output embedding, it is already a vector of size $L$ that can be analyzed by the MLP. For intermediate activations to be analyzed (convolutional layers), they need to be transformed into a vector. Based on existing literature [35, 53], data used for training maximizes the activations of intermediate neurons. Therefore, it seems reasonable to use the maximum value for each channel, obtaining $C$ values for each group of activations. We have also explored other options beyond channel maximum were explored (e.g., mean value), but they yielded inferior results and thus were discarded. This way, we transform the activations from $H\times W\times C$ into a vector of size $C$, which can be analyzed by a MLP. If the owner of the model provides access to this information, activations and embeddings can be analyzed together or separately, depending on what yields better results. More details regarding the specific architecture of the Vanilla MINT Model are provided in Fig. 4 (a).
@@ IV Experimental Framework
IV-A Face Recognition Data and Models
@@
As external data ($\mathcal{E}$) to train and test the MINT model $T$ we use the following datasets: The IJB-C [61] (3.5K identities and 138K face images), FDDB [62] (5.2K face images), GANDiffFace [63, 64] (10K identities and 500K face images), and Adience [65] (2.2K identities 26.5K face images).
IV-B Experimental Protocol
@@
  1. 1.
    No identity overlap between Train and Evaluation, ensuring that performance is not influenced by identity recognition rather than the primary task of recognizing training and non-training data.
  2. 2.
@@
  3. 3.
    In the evaluation, there is an equal number of samples for each class (the same from $\mathcal{D}$ and $\mathcal{E}$) to ensure balanced evaluation results. For example, as the Adience database $E$ used for evaluation in “Baseline Case” has 26.5K images, we took the same number of images from $D$, leaving the rest of the $D$ database available for training.
IV-B1 Data Pre-processing
@@
IV-B2 Auxiliary Auditable Data and Model Outcome
@@
IV-B3 MINT Models Training
With the previously obtained AAD and Model Outcome, we propose two MINT models. Depending on how we handle the acquired information, we propose two architectures (Sec. III-C): a Vanilla MLP model and a CNN model. The CNN utilizes the activations of the convolutional blocks of $M$ as-is, while the Vanilla model, as previously explained, applies max pooling per channel to convert the activations of the convolutional blocks into a vector that can be processed by the MLP.

### v3

@@ Abstract
journal: Information Fusion
[…]
[1]organization=Biometrics and Data Pattern Analytics Lab, Universidad Autonoma de Madrid, country=Spain
1 Introduction
In June 2023, the European Parliament adopted its position with respect to Artificial Intelligence (AI) [1]. Concretely, they have imposed some obligations to all AI companies to guarantee the protection of the citizen rights. For example, this new regulation imposes the registration of AI/ML models in a European Union (EU) database and gives national authorities of EU countries the power to request access to both the trained AI models and the associated procedural details conducted in the development of those models. This new regulation is a game changer, imposing more transparency in the development of AI technologies in Europe. As a result, novel auditing tools based on the access to trained AI models must be developed to monitor AI technologies and their correct deployment in our society. This trend is not limited to the European level, as the White House [2] has also taken a stance by publishing a memorandum that declares AI a matter of national security, emphasizing the need for access to AI technologies to review them and ensure the protection of citizens’ rights.
@@ 1 Introduction
This objective is related to the field of Membership Inference Attacks (MIAs), privacy attacks that target AI models trained on sensitive data [15, 16]. MIAs seeks to infer membership information about the training dataset used by exploiting the behavior and responses of the model. Also, MIAs leverage the potential memorization of the training data used by the AI models, especially when the models overfit [17, 18] or when the training data contain unique or distinctive patterns [15]. However, recent studies question MIAs effectiveness due to the inherent complexity of the task [19, 20].
@@
While MIAs are traditionally considered as attacks to quantify the potential privacy leakage of AI models, in the present article we introduce a novel perspective in line with the new AI regulation [1, 2]. Concretely, we consider it as an auditing tool for detecting the potential use of unauthorized data for training AI models, a concept we term as Membership Inference Test (MINT). The aim of MIAs and MINT is similar: to determine whether a given data sample was used during the training of a target AI model. However, although MIAs and our proposed MINT share the same objective, they diverge in their environmental conditions, leading to different methodological approaches. In MIAs, the model is trained assuming the role of an attacker who does not have access to the original model, but to a replica as similar as possible to it (shadow model) [15]. In MINT, we assume the role of an auditor who has direct access to the original model or partial information of it (e.g., intermediate activations). This is aligned with the new AI regulation from USA [2] and EU [1], which imposes the registration of the trained AI models in an EU database. As a result, this new paradigm moves from the attacker to the auditor point of view, allowing the development of MINT models under different environmental conditions.
@@
In the present study we concentrate on the challenging task of Face Recognition, due to the sensitive nature of face biometrics data [21, 22, 23] according to the General Data Protection Regulation (GDPR) promoted by EU [24], the California Consumer Privacy Act (CCPA) [25], and the Biometric Information Privacy Act (BIPA) [26]. Although MIAs and MINT differ due to their distinct environmental conditions, they also share significant similarities, so methods in one domain can inspire advancements in the other as well. The application of MIAs in Face Recognition remains underexplored, with most studies focusing on simple models featuring limited convolutional layers and designed for classification [27, 28, 29, 30]. These works do not address state-of-the-art models used in Face Recognition, leaving a significant gap in the literature. This gap in the state of the art is particularly crucial as MIAs frequently deal with classification tasks, using the network’s output vector to determine membership in the training data, like a form of ‘probability’ of belonging to each class [20, 31, 32, 33, 17]. However, due to the characteristics of Face Recognition systems, focused on recognition rather than classification, such vector usually does not exist. This multiclass probability vector is traditionally replaced by an embedding feature vector representing individual characteristics in a learned space. Consequently, the methods and results diverge even more from what we typically see in state-of-the-art MIAs studies [20].
Some recent studies have explored the application of Membership Inference Attacks (MIAs) on realistic Face Recognition models ([34, 35, 36]). However, these works focus on determining whether an entire user was included in the training set, leveraging specific model characteristics. For instance, [34, 35] demonstrated that users present in the training set tend to have their embedding feature vectors closer to each other. Similarly, [36] utilized guided captioning on CLIP models to detect the presence of users in the training set.
The detection of users in a model’s training set, while interesting, has notable limitations compared to the detection of specific samples used in training, which is the focus of MINT. User detection cannot pinpoint which specific samples were utilized during training; for instance, even if consent is provided for certain images, it does not mean all images of a person can be used. Sample detection, as addressed in this work, resolves this issue by identifying precisely which samples were included. Moreover, sample detection inherently achieves user detection, but the reverse does not hold, making the approach in this work more precise and granular. Additionally, user detection methods are restricted to face recognition models, whereas the methodologies proposed in this work are generalizable to any type of model. In fact, similar ideas are applied to other types of models in Appendix A.
@@
  • 1.
    We introduce MINT as a novel perspective to detect if given data was used in an AI model’s training from an AI auditor standpoint.
  • 2.
@@
The article is structured as follows. Sec. 2 provides a revision of the state of the art in the scope of the article. Sec. 3 describes the main concepts of the proposed MINT, and the specific details for the task considered, Face Recognition. The databases, Face Recognition systems, and experimental protocol are described in Sec. 4. Results are discussed in Sec. 5, whereas the significance of Face Recognition in MINT and the challenges of developing this technology in real-world applications are discussed in Sec. 6 and 7. Discussion and conclusions are finally drawn in Sec. 8 and 9.
2 Related Works
First, it is important to remark that, to the best of our knowledge, the specific problem addressed in the present article (MINT) has not been directly tackled in the current literature. Nevertheless, there are studies and research directions that touch upon similar topics or are grounded in analogous principles, which we will briefly discuss in this section. Our study closely relates to three fundamental research directions: “Membership Inference Attacks”, “Reconstructing Training Data from Neural Networks”, and “Data Tracing”. In the following, we provide concise explanations of these three primary research areas, with special emphasis on MIAs due to its significant similarity to our own research.
2.1 Membership Inference Attacks
The capability to detect the data used for training AI models presents a security concern that can potentially unveil sensitive [3] or private information [37], encompassing areas such as shopping preferences, health records, and pictures, among many others. This capability has been explored in the literature mainly under the name of Membership Inference Attacks (MIAs). These attacks aim to extract such sensitive information through an adversarial approach, without access to the model or its training data, as providers are hesitant to disclose such data. The pioneering work by Shokri et al. [15] laid the groundwork for this line of research. In their approach, they used “shadow models” that emulate the functionality of the original model, despite having no direct access to it. To construct these so-called “shadow models”, it is essential to have knowledge about the architecture of the original model they are based on, along with a subset of samples and statistics derived from the original dataset used for training. As a result, they crafted “shadow training sets” mirroring the original dataset, subsequently training shadow models to mimic the original model. This approach offers complete control over the training and non-training data of these models. A binary classifier was trained to distinguish between the training and non-training data used in the shadow models. Images were passed through the network, and output embeddings from all shadow models were used to train a simple binary classifier. Better results were observed with an increasing number of shadow models.
@@ 2.1 Membership Inference Attacks
More recently, the authors of [19] delved into the challenges of MIAs and the inherent complexity of the task. They concluded that the actual performance in the task is notably worse than the results published in the state of the art. This is largely due to the insufficient evaluation methodology. Furthermore, they explored previously untapped White-box information involving gradients, although they were not able to significantly improve the Black-box framework performance. Shafran et al. [42] analyzed how factors such as task complexity and image resolution impact the results of MINT.
2.2 Reconstructing Training Data from AI Models
@@ 2.3 Data Tracing
3 Membership Inference Test (MINT): Proposed Approach
MINT is a Membership Inference Test for a neural network’s training data that aims to comply with current legislation [1, 2], hence, assuming a certain level of collaboration from the model developer. The ultimate goal is to apply the MINT techniques to all data types (text, images, audio, etc.), considering varying levels of information provided by the model developer about the model training and operation. This spans from scenarios where no data information is provided (unsupervised) or very little (semi-supervised), up to when sufficient information about the training data is granted (supervised). Simultaneously, it considers whether you have access to the entire model (white-box) or merely to the model’s output (black-box). In the present study, we analyze a broad range of scenarios, varying the quantity of data available for training the MINT model and contemplating both white-box and black-box settings.
Detecting the data used to train a model is a complex task in some sense opposite to the primary goal of common machine learning (ML) problems: generalize well during inference to samples unseen during learning, i.e., behaving similarly for both training and non-training samples. Consequently, in standard ML we expect samples of the same class to be proximate in the model representation space, regardless of their presence in the training set (Fig. 2). MIAs and MINT on the other hand, seek to distinguish samples (used or not for training the Audited Model $M$) that the model $M$ intends to represent as close as possible. Hence, much of the existing MIA literature revolves around overfitted networks that do not generalize well [15, 55, 56, 20, 57, 58]. In this work, however, we apply MINT to state-of-the-art FR models, rather than overfitted models, to evaluate the real applicability of our method.
In MINT, we assume the role of model auditors with access to the original model, or at least to partial information of it. This is a consistent scenario either due to the legal considerations outlined in the present study (i.e., granting national authorities the possibility to request the model [1, 2]), or as a result of developer transparency initiatives aimed at end-users. This is an important distinction from related works around MIAs, as here we do not need to train “shadow models” to simulate the original model’s behavior. The “shadow models”, while a useful approximation, operate in a space different from the original (Fig. 2 left), potentially imposing limitations on the achievable performance [19]. Instead, with access to the original model, we can directly apply our proposed approach, which may yield different detection results and methods for similar tasks.
@@ 3 Membership Inference Test (MINT): Proposed Approach
3.1 Problem Statement and Terminology
Let us consider a Training Dataset ($\mathcal{D}$), an External Dataset ($\mathcal{E}$) and a collection of samples ($d\in\mathcal{D}\cup\mathcal{E}$). We assume a learned model ($M$) that is trained for a specific task (e.g., text generation, Face Recognition, etc.) using the dataset $\mathcal{D}$. For any input data record ($d$), the model ($M$) generates an outcome ($y$) based on $d$ and a set of parameters (w) learned during the training process, i.e., $y=M(d|\textbf{w})$.
@@ 3.1 Problem Statement and Terminology
  • 3.
    External Data $\mathcal{E}$: any data out of the collection $\mathcal{D}$.
  • 4.
@@
  • 5.
    Auxiliary Auditable Data $\textrm{AAD}=N(d|\textbf{w}^{\prime})$: intermediate outcomes of $M$ that result from processing an input data $d$ using a subset $\textbf{w}^{\prime}$ of the parameters w. The model outcome can be seen as the case where $^{\prime}=\textbf{w}$ w in which case $N(d|\textbf{w}^{\prime})=M(d|\textbf{w})$.
  • 6.
@@
3.2 MINT: Application to Face Recognition Models
In the previous Sec. 3.1, we have introduced the general concepts of our MINT proposal, which can be applied to any type of Audited Model $M$. Next, we detail the various components of the proposal for the application of MINT to the challenging task of Face Recognition.
@@ 3.2 MINT: Application to Face Recognition Models
3.3 Proposed MINT Models
@@
  • 1.
    Vanilla MINT Model: We propose a MLP network to capture the training patterns within the vectorized information of the audited model ($M$). When it comes to the output embedding, it is already a vector of size $L$ that can be analyzed by the MLP. For intermediate activations to be analyzed (convolutional layers), they need to be transformed into a vector. Based on existing literature [43, 61], data used for training maximizes the activations of intermediate neurons. Therefore, it seems reasonable to use the maximum value for each channel, obtaining $C$ values for each group of activations. Other alternatives (e.g., channel mean) were tested, but they yielded inferior results and thus were discarded. This way, we transform the activations from $H\times W\times C$ into a vector of size $C$, which can be analyzed by a MLP. If the owner of the model provides access to this information, activations and embeddings can be analyzed together or separately, depending on what yields better results. More details regarding the specific architecture of the Vanilla MINT Model are provided in Fig. 4 (a).
  • 2.
@@ 4 Experimental Framework
4.1 Face Recognition Data and Models
@@
As external data ($\mathcal{E}$) to train and test the MINT model $T$ we use the following datasets: The IJB-C [69] (3.5K identities and 138K face images), FDDB [70] (5.2K face images), GANDiffFace [71, 72] (10K identities and 500K face images), and Adience [73] (2.2K identities 26.5K face images). As mentioned in Sec. 3.2, it is crucial to avoid overlap between $\mathcal{E}$ and $\mathcal{D}$, as this would lead to incorrect labels, compromising the training of MINT Models $T$. As noted by Schlett et al. [74], there is a significant number of duplicates across commonly used datasets. Although the data in $\mathcal{E}$ is sourced from evaluation datasets that strive to prevent such overlap, the necessary tests have been conducted to ensure that this issue does not occur.
4.2 Experimental Protocol
@@
  • 1.
    No identity overlap when dividing the data between Train and Evaluation, ensuring that performance is not influenced by identity recognition rather than the primary task of recognizing training and non-training data.
  • 2.
@@
  • 3.
    In the evaluation, there is an equal number of samples for each class (the same from $\mathcal{D}$ and $\mathcal{E}$) to ensure balanced evaluation results. As the Adience database $E$ used for evaluation in “Baseline Case” has 26.5K images, we took the same number of images from $D$, leaving the rest of the $D$ database available for training.
4.2.1 Data Pre-processing
@@
4.2.2 Auxiliary Auditable Data and Model Outcome
@@
4.2.3 MINT Models Training
With the previously obtained AAD and Model Outcome, we propose two MINT models. Depending on how we handle the acquired information, we propose two architectures (Sec. 3.3): a Vanilla MLP model and a CNN model. The CNN utilizes the activations of the convolutional blocks of $M$ as-is, while the Vanilla model, as previously explained, applies max pooling per channel to convert the activations of the convolutional blocks into a vector that can be processed by the MLP.